OpenStudy (anonymous):

find the limit x approaches infinity, x-1/2x^2+3

OpenStudy (anonymous):

i used the direct substitution and i got \[-\frac{ 1 }{ 3 }\] for the right side and \[\frac{ 1 }{ 3}\] for the left side. Is that right?

OpenStudy (anonymous):

sorry, it is x approaches infinity.

OpenStudy (unklerhaukus):

oh

OpenStudy (unklerhaukus):

can you use l'Hopital's rule ?

OpenStudy (unklerhaukus):

divide the numerator and denominator by x \[\lim_{x\to\infty}\frac{x-1}{2x^2+3}\leadsto\frac\infty\infty\\ \lim_{x\to\infty}\frac{(x-1)/x}{(2x^2+3)/x}\]

OpenStudy (anonymous):

@UnkleRhaukus it tends to zero

OpenStudy (unklerhaukus):

but we haven't explained to get there yet

OpenStudy (anonymous):

you should tell him that the changes in denom are much greater than in nume then the whole of expression will approach zero..

OpenStudy (unklerhaukus):

yeah that's a different method , the powers of x in the denominator are greater than the powers of x in the numberator

OpenStudy (unklerhaukus):

\[\lim_{x\to\infty}\frac{x-1}{2x^2+3}\\=\lim_{x\to\infty}\frac{O(x)}{O(x^2)}\\=\lim_{x\to\infty}O(x^{-1})\]

OpenStudy (anonymous):

yeah, actually and it would be zero as x goes to infinity..

OpenStudy (unklerhaukus):

I like l'Hopital's relue method best \[\lim_{x\to\infty}\frac{x-1}{2x^2+3}\leadsto\frac\infty\infty\\ \stackrel{\text{l'H}}=\lim_{x\to\infty}\frac{\frac{\mathrm d}{dx}(x-1)}{\frac{\mathrm d}{dx}(2x^2+3)}\\=\lim_{x\to\infty}\frac{1}{4x}\\=\]

OpenStudy (unklerhaukus):

*rule

OpenStudy (anonymous):

it is a bad theory

OpenStudy (unklerhaukus):

bad? how so?

OpenStudy (anonymous):

it kills your creativity.

OpenStudy (unklerhaukus):

didn't kill mine

OpenStudy (unklerhaukus):

it doesn't even work every time

OpenStudy (rational):

lol L'hops is to limits what FTC is to riemann sums. good to use L'hops when you're allowed to

OpenStudy (anonymous):

it is okay only when you are going to make sure of what you've got by direct calculations..

OpenStudy (rational):

taylor series of top and bottom would be more creative (save degrees of a fraction logic)

OpenStudy (ikram002p):

well sincethe greates power of numerator is less than the greatest power of denominator i would say limit is zero ( without thinking ) hmmm idk if i have to think on this xD

OpenStudy (rational):

that was already said 2-3 times in above replies :P

OpenStudy (anonymous):

actually ...

OpenStudy (anonymous):

the problem is that how @dzie could ever get some nonzero answers.

OpenStudy (anonymous):

thank you all.

OpenStudy (anonymous):

but does that mean this cannot be solved using direct substitution?

OpenStudy (anonymous):

i need more explanation please on why we need to divide the numerator and the denominator by x or why we need to use L'hopital rule

OpenStudy (unklerhaukus):

direct substitution leads to infinity / infinity, which is an indeterminate form, so some technique must be applied to simplify the expression before the substitution is made

OpenStudy (anonymous):

but why is it \[\frac{ \infty }{ \infty }\] this was what i did \[\frac{ \infty -1 }{ 2(\infty)^{2}+3}\]

OpenStudy (unklerhaukus):

\[\infty-1=\infty\\ 2\times\infty^2+3=\infty\]

OpenStudy (anonymous):

okay. thanks